Find parametric equation of x=a(1+cos theta)
1
Expert's answer
2021-10-13T17:44:51-0400

Given polar form of equation of cardioid


"r=a(1+\cos \theta)"

We have that


"x=r\cos \theta"

"y=r\sin \theta"


Substitute


"x=a\cos \theta(1+\cos \theta)"

"y=a\sin \theta(1+\cos \theta)"

Replace "\theta" with "t" . The parametric equation is


"\begin{matrix}
x=a\cos t(1+\cos t)\\
y=a\sin t(1+\cos t)
\end{matrix}"